# Lost in Sequence: Do Large Language Models Understand Sequential Recommendation?

### v3

@@ 6. Conclusion
Appendix A Ethics Statement
To the best of our knowledge, this paper aligns with the KDD Code of Ethics without any ethical concerns. The datasets and codes employed in our research are publicly available.
Appendix B Details of LLM4Rec Prompt Construction
@@
B.1. Next Item Title Generation
@@
B.2. Next Item Retrieval
@@ Appendix D Baselines
Appendix E Implementation Details
@@
Appendix F Additional Experiments
F.1. Shuffled Training: Test Performance Curve
Figure 7 shows the test performance curves for each model during training.
F.2. Effectiveness of Input Prompts
@@
| NDCG@20 | 0.3405 | 0.3424 | |
| HR@10 | 0.4725 | 0.4885 | |
| HR@20 | 0.6239 | 0.6385 |
F.3. Distillation with Contrastive Learning
@@
| Distillation Loss | Movies | Scientific | Electronics | ||||||||||
| NDCG@10 | NDCG@20 | HR@10 | HR@20 | NDCG@10 | NDCG@20 | HR@10 | HR@20 | NDCG@10 | NDCG@20 | HR@10 | HR@20 | ||
| Contrastive | Original | 0.3410 | 0.3749 | 0.5345 | 0.6687 | 0.2767 | 0.3152 | 0.4817 | 0.6338 | 0.2553 | 0.2935 | 0.4277 | 0.5792 |
| Shuffle | 0.3151 | 0.3480 | 0.4975 | 0.6326 | 0.2638 | 0.3021 | 0.4650 | 0.6177 | 0.2398 | 0.2785 | 0.4065 | 0.5608 | |
| Change ratio | (-7.60%) | (-7.18%) | (-6.92%) | (-5.40%) | (-4.66%) | (-4.16%) | (-3.47%) | (-2.54%) | (-6.07%) | (-5.11%) | (-4.96%) | (-3.18%) | |
| LLM-SRec (MSE) | Original | 0.3560 | 0.3924 | 0.5569 | 0.7010 | 0.3388 | 0.3758 | 0.5532 | 0.6992 | 0.3044 | 0.3424 | 0.4885 | 0.6385 |
| Shuffle | 0.3272 | 0.3631 | 0.5169 | 0.6592 | 0.3232 | 0.3605 | 0.5336 | 0.6813 | 0.2845 | 0.3234 | 0.4638 | 0.6184 | |
| Change ratio | (-8.10%) | (-7.47%) | (-7.18%) | (-5.96%) | (-4.60%) | (-4.07%) | (-3.54%) | (-2.56%) | (-6.53%) | (-5.55%) | (-5.06%) | (-3.15%) | |
F.4. Preventing Over-smoothing
To validate that the prevents the over-smoothing problem, we measured the pairwise Euclidean distance between all user representations with and without the application of the under the four datasets. As shown in Table 13, applying the , i.e. LLM-SRec, results in larger distances between users than absence of , i.e., w.o. , which indicates that it helps generate more distinct representations for each user and consequently prevents the over-smoothing problem as described in Sec. 3.2.
| Movies | Scientific | Electronics | CDs | |
| w.o. | 7.69 | 7.78 | 8.03 | 9.82 |
| LLM-SRec | 9.33 | 11.54 | 13.81 | 11.68 |
F.5. Auto-regressive Training
@@
F.6. Additional Case Study
In Figure 9, we present an additional case study of what is shown in Sec. 4.3.3, which shows similar results. In case (a), the user’s preference shifts from speaker products to home security products made by a brand called ”Ring”. LLM-SRec correctly recommends a ”Ring” brand home security product, while SASRec captures the preference change but fails to make the correct recommendation. A-LLMRec, failing to leverage sequential information, still recommends a speaker product and thus fails to make the correct recommendation. This again highlights the importance of utilizing both sequential and textual information for accurate recommendations. In case (b), the user’s preference focuses on ”GoPro” brand products. Both LLM-SRec and A-LLMRec successfully recommend ”GoPro” products by leveraging textual knowledge. This emphasizes the crucial role of textual information in generating accurate recommendations. In case (c), the user’s preference shifts from ”Apple” brand products to ”SanDisk” products. LLM-SRec and SASRec successfully capture this changing preference, while A-LLMRec fails to capture the shift and makes an incorrect recommendation. This underscores the importance of capturing sequential information for accurately predicting user preferences. These cases demonstrate that leveraging both sequential and textual information is essential for making accurate recommendations.

### v4

@@ 6. Conclusion
Acknowledgements.
This work was supported by NAVER Corporation, the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00335098), and the National Research Foundation of Korea(NRF) funded by Ministry of Science and ICT (RS-2022-NR068758).References
- (1)
- AI@Meta (2024) AI@Meta. 2024. Llama 3 Model Card. (2024). https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md
- Bao et al. (2023) Keqin Bao, Jizhi Zhang, Yang Zhang, Wenjie Wang, Fuli Feng, and Xiangnan He. 2023. Tallrec: An effective and efficient tuning framework to align large language model with recommendation. In Proceedings of the 17th ACM Conference on Recommender Systems. 1007–1014.
- Chaney et al. (2015) Allison JB Chaney, David M Blei, and Tina Eliassi-Rad. 2015. A probabilistic model for using social networks in personalized item recommendation. In Proceedings of the 9th ACM Conference on Recommender Systems. 43–50.
- Choi et al. (2025) Seungyoon Choi, Sein Kim, Hongseok Kang, Wonjoong Kim, and Chanyoung Park. 2025. Dynamic Time-aware Continual User Representation Learning. arXiv preprint arXiv:2504.16501 (2025).
- Dai et al. (2023) Sunhao Dai, Ninglu Shao, Haiyuan Zhao, Weijie Yu, Zihua Si, Chen Xu, Zhongxiang Sun, Xiao Zhang, and Jun Xu. 2023. Uncovering chatgpt’s capabilities in recommender systems. In Proceedings of the 17th ACM Conference on Recommender Systems. 1126–1132.
- Geng et al. (2022) Shijie Geng, Shuchang Liu, Zuohui Fu, Yingqiang Ge, and Yongfeng Zhang. 2022. Recommendation as language processing (rlp): A unified pretrain, personalized prompt & predict paradigm (p5). In Proceedings of the 16th ACM Conference on Recommender Systems. 299–315.
- Harte et al. (2023) Jesse Harte, Wouter Zorgdrager, Panos Louridas, Asterios Katsifodimos, Dietmar Jannach, and Marios Fragkoulis. 2023. Leveraging large language models for sequential recommendation. In Proceedings of the 17th ACM Conference on Recommender Systems. 1096–1102.
- He et al. (2017) Xiangnan He, Lizi Liao, Hanwang Zhang, Liqiang Nie, Xia Hu, and Tat-Seng Chua. 2017. Neural collaborative filtering. In Proceedings of the 26th international conference on world wide web. 173–182.
- Hidasi et al. (2015) Balázs Hidasi, Alexandros Karatzoglou, Linas Baltrunas, and Domonkos Tikk. 2015. Session-based recommendations with recurrent neural networks. arXiv preprint arXiv:1511.06939 (2015).
- Hou et al. (2024a) Yupeng Hou, Jiacheng Li, Zhankui He, An Yan, Xiusi Chen, and Julian McAuley. 2024a. Bridging Language and Items for Retrieval and Recommendation. arXiv preprint arXiv:2403.03952 (2024).
- Hou et al. (2024b) Yupeng Hou, Junjie Zhang, Zihan Lin, Hongyu Lu, Ruobing Xie, Julian McAuley, and Wayne Xin Zhao. 2024b. Large language models are zero-shot rankers for recommender systems. In European Conference on Information Retrieval. Springer, 364–381.
- Hu et al. (2022) Edward J Hu, yelong shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2022. LoRA: Low-Rank Adaptation of Large Language Models. In International Conference on Learning Representations.
- Kang and McAuley (2018) Wang-Cheng Kang and Julian McAuley. 2018. Self-attentive sequential recommendation. In 2018 IEEE international conference on data mining (ICDM). IEEE, 197–206.
- Kaplan et al. (2020) Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. 2020. Scaling laws for neural language models. arXiv preprint arXiv:2001.08361 (2020).
- Kim et al. (2025) Jiwan Kim, Hongseok Kang, Sein Kim, Kibum Kim, and Chanyoung Park. 2025. Disentangling and Generating Modalities for Recommendation in Missing Modality Scenarios. arXiv preprint arXiv:2504.16352 (2025).
- Kim et al. (2024) Sein Kim, Hongseok Kang, Seungyoon Choi, Donghyun Kim, Minchul Yang, and Chanyoung Park. 2024. Large Language Models meet Collaborative Filtering: An Efficient All-round LLM-based Recommender System. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (Barcelona, Spain) (KDD ’24). Association for Computing Machinery, New York, NY, USA, 1395–1406. doi:10.1145/3637528.3671931
- Kim et al. (2023) Sein Kim, Namkyeong Lee, Donghyun Kim, Minchul Yang, and Chanyoung Park. 2023. Task Relation-aware Continual User Representation Learning. In Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 1107–1119.
- Klenitskiy et al. (2024) Anton Klenitskiy, Anna Volodkevich, Anton Pembek, and Alexey Vasilev. 2024. Does It Look Sequential? An Analysis of Datasets for Evaluation of Sequential Recommendations. In Proceedings of the 18th ACM Conference on Recommender Systems (Bari, Italy) (RecSys ’24). Association for Computing Machinery, New York, NY, USA, 1067–1072. doi:10.1145/3640457.3688195
- Krizhevsky et al. (2012) Alex Krizhevsky, Ilya Sutskever, and Geoffrey E Hinton. 2012. Imagenet classification with deep convolutional neural networks. Advances in neural information processing systems 25 (2012).
- Li et al. (2023c) Jiacheng Li, Ming Wang, Jin Li, Jinmiao Fu, Xin Shen, Jingbo Shang, and Julian McAuley. 2023c. Text is all you need: Learning language representations for sequential recommendation. In Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 1258–1267.
- Li et al. (2023a) Xiangyang Li, Bo Chen, Lu Hou, and Ruiming Tang. 2023a. Ctrl: Connect tabular and language model for ctr prediction. CoRR (2023).
- Li et al. (2023b) Xinhang Li, Chong Chen, Xiangyu Zhao, Yong Zhang, and Chunxiao Xing. 2023b. E4srec: An elegant effective efficient extensible solution of large language models for sequential recommendation. arXiv preprint arXiv:2312.02443 (2023).
- Liao et al. (2024) Jiayi Liao, Sihang Li, Zhengyi Yang, Jiancan Wu, Yancheng Yuan, Xiang Wang, and Xiangnan He. 2024. LLaRA: Large Language-Recommendation Assistant. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (Washington DC, USA) (SIGIR ’24). Association for Computing Machinery, New York, NY, USA, 1785–1795. doi:10.1145/3626772.3657690
- Liu and Abbeel (2024) Hao Liu and Pieter Abbeel. 2024. Blockwise parallel transformers for large context models. Advances in Neural Information Processing Systems 36 (2024).
- Liu (2019) Yinhan Liu. 2019. Roberta: A robustly optimized bert pretraining approach. arXiv preprint arXiv:1907.11692 364 (2019).
- Mnih and Salakhutdinov (2007) Andriy Mnih and Russ R Salakhutdinov. 2007. Probabilistic matrix factorization. Advances in neural information processing systems 20 (2007).
- Oh et al. (2023) Yunhak Oh, Sukwon Yun, Dongmin Hyun, Sein Kim, and Chanyoung Park. 2023. Muse: music recommender system with shuffle play recommendation enhancement. In Proceedings of the 32nd ACM international conference on information and knowledge management. 1928–1938.
- Rendle et al. (2010) Steffen Rendle, Christoph Freudenthaler, and Lars Schmidt-Thieme. 2010. Factorizing personalized markov chains for next-basket recommendation. In Proceedings of the 19th international conference on World wide web. 811–820.
- Sanner et al. (2023) Scott Sanner, Krisztian Balog, Filip Radlinski, Ben Wedin, and Lucas Dixon. 2023. Large language models are competitive near cold-start recommenders for language-and item-based preferences. In Proceedings of the 17th ACM conference on recommender systems. 890–896.
- Sun et al. (2019) Fei Sun, Jun Liu, Jian Wu, Changhua Pei, Xiao Lin, Wenwu Ou, and Peng Jiang. 2019. BERT4Rec: Sequential recommendation with bidirectional encoder representations from transformer. In Proceedings of the 28th ACM international conference on information and knowledge management. 1441–1450.
- Takahagi and Shinnou (2023) Kyosuke Takahagi and Hiroyuki Shinnou. 2023. Data Augmentation by Shuffling Phrases in Recognizing Textual Entailment. In Proceedings of the 37th Pacific Asia Conference on Language, Information and Computation. Association for Computational Linguistics, Hong Kong, China, 194–200. https://aclanthology.org/2023.paclic-1.19/
- Tang and Wang (2018a) Jiaxi Tang and Ke Wang. 2018a. Personalized Top-N Sequential Recommendation via Convolutional Sequence Embedding. In Proceedings of the Eleventh ACM International Conference on Web Search and Data Mining (Marina Del Rey, CA, USA) (WSDM ’18). Association for Computing Machinery, New York, NY, USA, 565–573. https://doi.org/10.1145/3159652.3159656
- Tang and Wang (2018b) Jiaxi Tang and Ke Wang. 2018b. Personalized top-n sequential recommendation via convolutional sequence embedding. In Proceedings of the eleventh ACM international conference on web search and data mining. 565–573.
- Wang et al. (2024) Chen Wang, Liangwei Yang, Zhiwei Liu, Xiaolong Liu, Mingdai Yang, Yueqing Liang, and Philip S. Yu. 2024. Collaborative Alignment for Recommendation. In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (Boise, ID, USA) (CIKM ’24). Association for Computing Machinery, New York, NY, USA, 2315–2325. doi:10.1145/3627673.3679535
- Wang et al. (2022) Chenyang Wang, Yuanqing Yu, Weizhi Ma, Min Zhang, Chong Chen, Yiqun Liu, and Shaoping Ma. 2022. Towards Representation Alignment and Uniformity in Collaborative Filtering. In Proceedings of the 28th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (Washington DC, USA) (KDD ’22). Association for Computing Machinery, New York, NY, USA, 1816–1825. doi:10.1145/3534678.3539253
- Wang and Isola (2020) Tongzhou Wang and Phillip Isola. 2020. Understanding contrastive representation learning through alignment and uniformity on the hypersphere. In International conference on machine learning. PMLR, 9929–9939.
- Woolridge et al. (2021) Daniel Woolridge, Sean Wilner, and Madeleine Glick. 2021. Sequence or Pseudo-Sequence? An Analysis of Sequential Recommendation Datasets.. In Perspectives@ RecSys.
- Wu et al. (2024) Junda Wu, Cheng-Chun Chang, Tong Yu, Zhankui He, Jianing Wang, Yupeng Hou, and Julian McAuley. 2024. Coral: Collaborative retrieval-augmented large language models improve long-tail recommendation. In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining. 3391–3401.
- Wu et al. (2019) Shu Wu, Yuyuan Tang, Yanqiao Zhu, Liang Wang, Xing Xie, and Tieniu Tan. 2019. Session-based recommendation with graph neural networks. In Proceedings of the AAAI conference on artificial intelligence, Vol. 33. 346–353.
- Yuan et al. (2019) Fajie Yuan, Alexandros Karatzoglou, Ioannis Arapakis, Joemon M Jose, and Xiangnan He. 2019. A simple convolutional generative network for next item recommendation. In Proceedings of the twelfth ACM international conference on web search and data mining. 582–590.
- Yue et al. (2023) Zhenrui Yue, Sara Rabhi, Gabriel de Souza Pereira Moreira, Dong Wang, and Even Oldridge. 2023. LlamaRec: Two-stage recommendation using large language models for ranking. arXiv preprint arXiv:2311.02089 (2023).
- Zhang et al. (2023) Yang Zhang, Fuli Feng, Jizhi Zhang, Keqin Bao, Qifan Wang, and Xiangnan He. 2023. Collm: Integrating collaborative embeddings into large language models for recommendation. arXiv preprint arXiv:2310.19488 (2023).
Appendix A Details of LLM4Rec Prompt Construction
@@
A.1. Next Item Title Generation
@@
A.2. Next Item Retrieval
@@ Appendix C Baselines
Appendix D Implementation Details
@@
Appendix E Additional Experiments
E.1. Effectiveness of Input Prompts
@@
| HR@10 | 0.4725 | 0.4885 |
E.2. Distillation with Contrastive Learning
@@
| Distillation Loss | Inference | Movies | Scientific | Electronics |
| Contrastive | Original | 0.3410 | 0.2767 | 0.2553 |
| Shuffle | 0.3151 | 0.2638 | 0.2398 | |
| Change ratio | (-7.60%) | (-4.66%) | (-6.07%) | |
| LLM-SRec (MSE) | Original | 0.3560 | 0.3388 | 0.3044 |
| Shuffle | 0.3272 | 0.3232 | 0.2845 | |
| Change ratio | (-8.10%) | (-4.60%) | (-6.53%) |
E.3. Auto-regressive Training